OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2).

OpenStudy (anonymous):

if 2 lines r parallel their slope r equal 3x - y = -3 slope= 3 equation of the line whse slope= 3 n passing through the point (4, -2). y-y1=slope(x-x1)

OpenStudy (anonymous):

here slope=3 n (x1 y1)=(4 -2) proceed

OpenStudy (anonymous):

okay thank you :) i got it now